the perimeter of a rectangle is 68 ft. The lenght is 2 ft less than 3 times the width. Please help me find the lenght and width.
Answer by Mathtut(3670)

P=68........eq 1
L=3W-2......eq 2
P=2L+2W.....eq 3
:
take L's value and P's value from eq 1 and 2 and plug them into eq 3
:
68=2(3W-2)+2W
:
68=6W-4+2W
:
8W=72
:
W=9 feet
:
L=3(9)-2=25 feet